OpenStudy (anonymous):

the rate of a stock falls 20% in 1 year and 25% in another two. How manz percent did it fall in two years total?

OpenStudy (anonymous):

After the first year we would have the stock at an 80% of the initial value. Considering that, unlike normal fluctuation, the stock would fall constantly in the next two years, we would have a 12.5% fall each year, summing up the 25%. So we have First Year (a) \[4x/5=y\] and the Second Year (b) \[77.5y/100\] replacing that we get \[(77.5 * 4x/5 ) / 100= 62x\] \[100 - 62 = 38\] so the stock falls a total of 38% in two years, waiting to fall another 18 in the next year

OpenStudy (anonymous):

i will give you a medal because you gave me some brains! but i am still confused about the the part where you have to substitude something for y. where did you get 4x/5 ?

OpenStudy (anonymous):

never mind :D the page did was not totaly loaded up! THank you!